OpenStudy (anonymous):

Simplify (10x2- 8x - 1) + (4x2- 8x - 5).

OpenStudy (anonymous):

@driftracer305

Nnesha (nnesha):

combine like terms

OpenStudy (anonymous):

would the first one be 14x2? or 14x4?

OpenStudy (driftracer305):

ok...so open up the parenthesis since u dont have a negative....nd put like terms together

Nnesha (nnesha):

\[10x ^{2} + 4x ^{2} = ???\]

Nnesha (nnesha):

no you don't have to add exponent

OpenStudy (anonymous):

sooo I got 14x2-16x-6 :) is that right?

Nnesha (nnesha):

when you multiply numbers with variables then you should add exponents

OpenStudy (driftracer305):

correcto

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

Yay! thanks so much :)

Nnesha (nnesha):

yayayayyaay!!!!!!!!!!!! good job

Nnesha (nnesha):

my pleasure :) :)

OpenStudy (driftracer305):

sigh!!....that was a lot of work!! *sarcasm*

OpenStudy (anonymous):

:c

OpenStudy (driftracer305):

we proceed to next one

OpenStudy (anonymous):

Alrihgty!

OpenStudy (xxdragoonxx):

(10x2-8x-1)+(4x2-8x-5) (-8x+2*10x-1)+(2*4x-8x-5) = (-8x+2*10x-1) = (20x-8x-1) = 12x-1 = 12x-1+(2*4x-8x-5) = (2*4x-8x-5) = (8x-8x-5) = (0-5) = (-5) = -5 = 12x-1-5 = 12x-6

Nnesha (nnesha):

@xxDragoonxx 2 is exponent of x not the multiply by 10 or 4